# Information Limits of Joint Community Detection and Finite Group Synchronization

### v1

@@ 5.1 Preliminaries
Figure 5: Simple examples on $E_{2,S_{1}}$. We consider five different scenarios: $(a)$ and $(b)$: no edge or only one edge connected between $S_{11}$ and $S_{12}$, then $E_{2,S_{1}}$ occurs with probability 1; $(c)$: assuming the subgraph formed by $S_{11}$ is connected, and two edges connected to the same node in $S_{12}$ form a cycle, then $E_{2,S_{1}}$ occurs when the cycle consistency satisfies with probability $M^{-1}$; $(d)$ and $(e)$: more edges between $S_{11}$ and $S_{12}$ leads to more cycles and further lower probability of $E_{2,S_{1}}$. See text for details.
To this end, by definition $E_{2,S_{1}}$ occurs as long as there exists a hypothesis of the group elements $\bm{g}$ that satisfies the consistency $g_{ij}=g_{i}g_{j}^{-1}$ on each edge in $S_{1}$. However, determining the conditional probability $\mathcal{P}_{\bm{x}^{*}}\left(E_{2,S_{1}}\mid r_{1}\right)$ appears to be challenging as it involves two sources of randomness:
@@
    For any pair of node $(i,j)$ within $S_{11}$ or $S_{12}$, they are connected with probability $p=\frac{a\log n}{n}$, and the group transformation observed is $g_{ij}=g_{i}^{*}(g_{j}^{*})^{-1}=0$ with probability 1
@@
    For any pair of nodes $(i,j)$ between $S_{11}$ and $S_{12}$, the edges are randomly assigned under the constraint that the total number of edges is $r_{1}$, and $g_{ij}$ is uniformly drawn from $\mathcal{G}_{M}$
Here, our analysis on $\mathcal{P}_{\bm{x}^{*}}\left(E_{2,S_{1}}\mid r_{1}\right)$ starts from several simple examples shown in Figure 5, which is illustrated as follows:
@@
[…]
    $g_{ij}=g_{i}^{*}(g_{j}^{*})^{-1}=0.$
    As a result, no matter how the nodes within $S_{11}$ or $S_{12}$ are connected, we have
    $\mathcal{P}_{\bm{x}^{*}}\left(E_{2,S_{1}}|r_{1}=0\right)=1$
    as a hypothesis $g_{i}=0,\;i=1,\ldots,n$ always satisfies the consistency.
@@
    $g_{i}=0,\;\forall i\in S_{11}\quad\text{and}\quad g_{i}=g_{l_{1}l_{2}},\;\forall i\in S_{12}$
[…]
@@
      The subgraph formed by $S_{11}$, denoted as $G(S_{11})$, is connected
@@
      The $r_{1}=2$ edges between $S_{11}$ and $S_{12}$ connect the same node $l_{1}\in S_{12}$
@@
    $l_{1}\;\xrightarrow{g_{l_{1}l_{2}}}\;l_{2}\;\xrightarrow{{\color[rgb]{1,1,1}g_{l_{1}l_{2}}}}\;\cdots\;\xrightarrow{{\color[rgb]{1,1,1}g_{l_{1}l_{2}}}}\;l_{3}\;\xrightarrow{g_{l_{3}l_{1}}}\;l_{1}$ (43)
[…]
    $g_{l_{1}l_{2}}\cdots g_{l_{3}l_{1}}=0$ (44)
@@
    $\mathcal{P}_{\bm{x}^{*}}\left(E_{2,S_{1}}|E_{c},r_{1}=2\right)=M^{-1}$
@@
    Now, we consider $r_{1}=3$ and assume an event $E_{d}$ similar to $E_{c}$ above occurs, i.e. $G(S_{11})$ is connected and all the three edges connect the same node $l_{1}\in S_{22}$. We denote the new node connect to $l_{1}$ as $l_{4}\in S_{11}$. As a result, this gives rise another cycle
    $l_{1}\;\xrightarrow{g_{l_{1}l_{4}}}\;l_{4}\;\xrightarrow{{\color[rgb]{1,1,1}g_{l_{1}l_{2}}}}\;\cdots\;\xrightarrow{{\color[rgb]{1,1,1}g_{l_{1}l_{2}}}}\;l_{3}\;\xrightarrow{g_{l_{3}l_{1}}}\;l_{1}$
[…]
    $\mathcal{P}_{\bm{x}^{*}}\left(E_{2,S_{1}}|E_{d},r_{1}=3\right)=M^{-2}$
@@
[…]
@@
    Notably, here we use $\leq$’ instead of ‘$=$’ since we have ignored the edge connections in $S_{12}$, and there might be more cycles exist by considering those edges. However, such an upper bound is sufficient to our analysis.
Now, our analysis in Figure 5 can be generalized to the following result
@@
Let $E_{\mathrm{con}}$ denotes the event that $G(S_{11})$ is connected. Then,
$\mathcal{P}_{\bm{x}^{*}}\left(E_{2,S_{1}}\mid r_{1},E_{\mathrm{con}}\right)\leq\min\left\{1,\;M^{\Delta-r_{1}}\right\}$ (45)
where $\Delta=|S_{12}|$.
@@
$\mathcal{P}_{\bm{x}^{*}}\left(E_{2,S_{1}}\mid r_{1},E_{\mathrm{con}}\right)$ $\leq\mathcal{P}\left(\bigcap_{i\in S_{12}}E_{\text{cycle},i}\right)\overset{(a)}{=}\prod_{i\in S_{12}}\mathcal{P}\left(E_{\text{cycle},i}\right)$
$=\prod_{ii\in S_{12}}M^{-\max\{0,\;r_{1,i}-1\}}=\prod_{i\in S_{12}}\min\left\{1,\;M^{1-r_{1,i}}\right\}$
@@
=$\overset{(b)}{=}\min\left\{1,\;M^{\Delta-r_{1}}\right\}$
@@
However, Lemma 5.1 is still one step away from the desired bound as it requires $G(S_{11})$ to be connected. Fortunately, as $G(S_{11})$ follows an Erdős–Rényi model and when $p=\Theta\left(\frac{\log n}{n}\right)$, we can show that at least a giant component of size
$(1-o(1))\cdot|S_{11}|$
exists in $G(S_{11})$ w.h.p. (see Theorem A.4). As a result, an upper bound similar to (45) can be obtained as follows:
@@
$\mathcal{P}_{\bm{x}^{*}}\left(E_{2,S_{1}}\mid r_{1}\right)\leq n^{-\Theta(\log n)}+\min\left\{1,M^{\Delta-(1-o(1))r_{1}}\right\}$ (46)
@@
Lemma 5.2 does not require any assumption on the graph and can be applied to (42) directly. The proof is deferred to Appendix B.1. Similarly,
@@
Plugging this into (42) yields
@@
Now, by plugging (47) into (39) and taking the summation over $r$ and $r^{*}$, we obtain the following bound on $\mathcal{P}_{\Delta}$. The details are left in Appendix B.2.
@@
$\mathcal{P}_{\Delta}\leq M^{2\Delta}\cdot\exp\left(-\frac{N_{\Delta}\log n}{n}\left(a+b-(2+o(1))\sqrt{\frac{ab}{M}}\right)\right)$
@@
Step 4: Sum over $\Delta$
@@
$\leq 2\sum_{\Delta=1}^{\left\lceil n/4\right\rceil}g(\Delta)$ (48)
where $(a)$ uses the property of binomial coefficient that $\binom{n}{k}\leq\left(\frac{n}{k}\right)^{k}$ and we define each term in the summation as $g(\Delta)$. To proceed, we take the logarithm of $g(\Delta)$ as
@@
where the last step uses $c\in\left[\frac{2c}{3},c\right]$. This completes the proof.
5.3 Proof of Theorem 3.2
Our proof technique is mainly inspired by the one of [1, Theorem 1]. Similar to the proof of Theorem 5.2, we still follow the assumption (32). To begin with, recall the MLE (27) fails to exactly recover the communities $\bm{\kappa}^{*}$ if and only if there exists an alternative hypothesis $\bm{x}=\{\bm{\kappa},\bm{g}\}$ that
@@
Case 1
We first consider when
$\frac{Mp(1-q)}{q(1-p)}\geq 1.$
By plugging in the scaling $p=\frac{a\log n}{n}$ and $q=\frac{b\log n}{n}$ this is equivalent to $\frac{Ma}{b}\geq 1$, and the MLE fails only if $r\geq r^{*}$.
For ease of presentation, without specification, we denote $\mathcal{P}(\cdot)$ as the probability conditioning on the ground truth $\bm{x}^{*}$ and omit the subscript. Then, let us define the following events:
@@
    node $i\in S_{1}^{*}$ (resp. $S_{2}^{*}$) is connecting to more nodes in $S_{2}^{*}$ (resp. $S_{1}^{*}$) than in $S_{1}^{*}$ (resp. $S_{2}^{*}$) i.e.
    $|\mathcal{E}(i,S_{1}^{*}\backslash i)|<|\mathcal{E}(i,S_{2}^{*})|$
    (resp. $|\mathcal{E}(i,S_{2}^{*}\backslash i)|<|\mathcal{E}(i,S_{1}^{*})|$)
@@
    node $i\in S_{1}^{*}$ (resp. $S_{2}^{*}$) satisfies $g_{ij}=g_{i}$, for any $(i,j)\in\mathcal{E}(i,S_{2}^{*})$ (resp. $\mathcal{E}(i,S_{1}^{*})$)
@@
By symmetry, we have $\mathcal{P}(E_{1})=\mathcal{P}(E_{2})\geq\frac{4}{5}$. Then, consider the event $E=E_{1}\cap E_{2}$ occurs, which indicates there exists a pair of nodes $i\in S_{1}^{*}$ and $j\in S_{2}^{*}$ with the two events $E^{(i)}$ and $E^{(j)}$ occur with some $g_{i}$ and $g_{j}$. As a result, one can construct an alternative hypothesis $\bm{x}$ from $\bm{x}^{*}$ by swapping the community assignment of $i$ and $j$ i.e. $i\in S_{2}$ and $j\in S_{1}$, and let the associated group elements be $g_{i}$ and $g_{j}$, then $\mathcal{P}(\bm{y}|\bm{x})>\mathcal{P}(\bm{y}|\bm{x}^{*})$, indicating $F$ occurs. Therefore,
$\mathcal{P}(F)\geq\mathcal{P}(E_{1}\cap E_{2})\geq\mathcal{P}(E_{1})+\mathcal{P}(E_{2})-1\geq\frac{3}{5}$
@@
$\gamma:=\log^{3}n,\quad\delta:=\frac{\log n}{\log\log n}$
and we assume them to be integers (otherwise round them to the nearest ones). Then, let $H$ be a fixed subset of $S_{1}^{*}$ of size $\frac{n}{\gamma}$, we further define the following two events
@@
    no node in $H$ connects to at least $\delta$ number of other nodes in $H$
@@
    node $i\in H$ satisfies $|\mathcal{E}(i,S_{1}^{*}\backslash H)|+\delta\leq|\mathcal{E}(i,S_{2}^{*})|$
@@
By definition, $E_{H}$ occurs when there exists a node $i\in H$ with some $g_{i}$ that such that $E_{H,A}^{(i)}$ and $E_{B}^{(i,g_{i})}$ occur. Then one can see that $F_{H}\cap E_{H}\subseteq E_{1}$, as $F_{H}\cap E_{H}^{(i)}$ indicates $E^{(i)}$ occurs. Therefore, by applying $\mathcal{P}(F_{H})\geq\frac{9}{10}$ from [1, Lemma 10], we have
@@
$\overset{(b)}{\geq}1-\prod_{i\in H}\left(1-\mathcal{P}\left(E_{H,A}^{(i)}\bigcap E_{B}^{(i,0)}\right)\right)$ (52)
@@
Here, $(a)$ comes from the fact that the two events $E_{H,A}^{(i)}$ and $E_{B}^{(i,0)}$ occurs independently given $r_{1}$ and $r_{2}$; $(b)$ holds as given $r_{1}$ and $r_{2}$, $E_{H,A}^{(i)}$ is deterministic, and by definition $E_{B}^{(i,0)}$ occurs only if $g_{ij}=0$ for all $(i,j)\in\mathcal{E}(i,S_{2}^{*})$, and each $g_{ij}=0$ occurs independent with probability $M^{-1}$. Following this way with some heavy calculation, we obtain
@@
Case 2.
Now we consider
$\frac{Mp(1-q)}{q(1-p)}<1$
or equivalently $\frac{Ma}{b}<1$, which indicates MLE fails only if $r\leq r^{*}$. The proof is very similar to case 1 where most of the result still applies at here. Therefore, we will skip those repeated analyses. To begin with, we reuse all the notations of events in case 1 but redefine the following two:

### v2

@@ 6.3 Step 3: Cycle consistency analysis
Figure 5: Some illustrations on $E_{2,S_{1}}$. We consider five different scenarios: (a) and (b): no edge or only one edge connected between $S_{11}$ and $S_{12}$, then $E_{2,S_{1}}$ occurs with probability 1; (c): assuming the subgraph formed by $S_{11}$ is connected, and two edges connected to the same node in $S_{12}$ form a cycle, then $E_{2,S_{1}}$ occurs when the cycle consistency is satisfied with probability $M^{-1}$; (d) and (e): more edges between $S_{11}$ and $S_{12}$ leads to more cycles and further lower probability of $E_{2,S_{1}}$. See text for details.
To this end, by definition, $E_{2,S_{1}}$ occurs as long as there exists a hypothesis of the group elements $\bm{g}$ that satisfies the consistency $g_{ij}=g_{i}g_{j}^{-1}$ on each edge in $S_{1}$. However, determining the conditional probability $\mathcal{P}_{\bm{x}^{*}}\left(E_{2,S_{1}}\mid r_{1}\right)$ appears to be challenging as it involves two sources of randomness:
@@
    For any pair of node $(i,j)$ within $S_{11}$ or $S_{12}$, they are connected with probability $p=\frac{a\log n}{n}$, and the group transformation observed is $g_{ij}=g_{i}^{*}(g_{j}^{*})^{-1}=I$ (recall the assumption (38) that $g_{i}^{*}=I,\;i=1,\ldots,n$) with probability 1.
@@
    For any pair of nodes $(i,j)$ between $S_{11}$ and $S_{12}$, the edges are randomly assigned under the constraint that the total number of edges between $S_{11}$ and $S_{12}$ is $r_{1}$ (recall the definition of $r_{1}$ in (46)), and $g_{ij}$ is uniformly drawn from $\mathcal{G}_{M}$.
Here, our analysis on $\mathcal{P}_{\bm{x}^{*}}\left(E_{2,S_{1}}\mid r_{1}\right)$ starts from several simple examples that correspond to different choices of $r_{1}$, as shown in Figure 5 that, which are elaborated as follows:
@@
    We start from Figure 5(a) with $r_{1}=0$ such that no edges connect between $S_{11}$ and $S_{12}$. In this case, $g_{ij}$ on all edges within $S_{11}$ and $S_{12}$ satisfy $g_{ij}=g_{i}^{*}(g_{j}^{*})^{-1}=I$. As a result, no matter how the nodes within $S_{11}$ or $S_{12}$ are connected, we have
    $\mathcal{P}_{\bm{x}^{*}}\left(E_{2,S_{1}}|r_{1}=0\right)=1,$
    as a hypothesis $g_{i}=I,\;i=1,\ldots,n$ always satisfies the consistency.
@@
    $g_{i}=I,\;\forall i\in S_{11}\quad\text{and}\quad g_{i}=g_{l_{1}l_{2}},\;\forall i\in S_{12},$
    as shown in Figure 5(b). Therefore, we still have
@@
      The subgraph formed by $S_{11}$, denoted as $G(S_{11})$, is connected.
@@
      The $r_{1}=2$ edges between $S_{11}$ and $S_{12}$ connect the same node $l_{1}\in S_{12}$.
@@
    $l_{1}\;\xrightarrow{g_{l_{1}l_{2}}}\;l_{2}\;\xrightarrow{{\color[rgb]{1,1,1}g_{l_{1}l_{2}}}}\;\cdots\;\xrightarrow{{\color[rgb]{1,1,1}g_{l_{1}l_{2}}}}\;l_{3}\;\xrightarrow{g_{l_{3}l_{1}}}\;l_{1},$ (52)
    which is shown in Figure 5(c). In this case, one can notice that $E_{2}$ occurs only if the cycle consistency is met, i.e.,
    $g_{l_{1}l_{2}}\cdots g_{l_{3}l_{1}}=I,$ (53)
@@
    $\mathcal{P}_{\bm{x}^{*}}\left(E_{2,S_{1}}|E_{c},r_{1}=2\right)=M^{-1},$
@@
    Now, we consider $r_{1}=3$ and assume an event $E_{d}$ similar to $E_{c}$ above occurs, i.e., $G(S_{11})$ is connected and all the three edges connect the same node $l_{1}\in S_{22}$. We denote the new node connect to $l_{1}$ as $l_{4}\in S_{11}$. As a result, this gives rise another cycle
    $l_{1}\;\xrightarrow{g_{l_{1}l_{4}}}\;l_{4}\;\xrightarrow{{\color[rgb]{1,1,1}g_{l_{1}l_{2}}}}\;\cdots\;\xrightarrow{{\color[rgb]{1,1,1}g_{l_{1}l_{2}}}}\;l_{3}\;\xrightarrow{g_{l_{3}l_{1}}}\;l_{1},$
    as shown in Figure 5(d). Again, by applying the same argument as in part $(c)$, one can see that the new cycle satisfies the cycle consistency with probability $M^{-1}$ and is independent with the old one. Therefore, we have
    $\mathcal{P}_{\bm{x}^{*}}\left(E_{2,S_{1}}|E_{d},r_{1}=3\right)=M^{-2},$
@@
    Lastly, we consider a more involved example when two nodes in $S_{12}$ connect with $S_{11}$, and the subgraph $G(S_{11})$ is still connected, as shown in Figure 5(e). We denote such an event as $E_{e}$. Here, each of the two nodes has three edge connections, and therefore there are at least four cycles. In this case, we have
@@
    Notably, here we use $\leq$’ instead of ‘$=$’ since we have ignored the edge connections in $S_{12}$, and there will be more cycles exist by considering those edges. However, such an upper bound is sufficient to our analysis.
Now, based on the previous discussions, our analysis in Figure 5 can be generalized to the following result
@@
Let $E_{\mathrm{con}}$ denotes the event that the subgraph $G(S_{11})$ is connected. Then,
$\mathcal{P}_{\bm{x}^{*}}\left(E_{2,S_{1}}\mid r_{1},E_{\mathrm{con}}\right)\leq\min\left\{1,\;M^{\Delta-r_{1}}\right\},$ (54)
where $\Delta:=|S_{12}|=|S_{21}|$ is defined in (40).
@@
$\mathcal{P}_{\bm{x}^{*}}\left(E_{2,S_{1}}\mid r_{1},E_{\mathrm{con}}\right)$ $\leq\mathcal{P}_{\bm{x}^{*}}\left(\bigcap_{i\in S_{12}}E_{\text{cycle},i}\mid r_{1},E_{\mathrm{con}}\right)\overset{(a)}{=}\prod_{i\in S_{12}}\mathcal{P}_{\bm{x}^{*}}\left(E_{\text{cycle},i}\mid r_{1},E_{\mathrm{con}}\right)$
$=\prod_{i\in S_{12}}M^{-\max\{0,\;r_{1,i}-1\}}=\prod_{i\in S_{12}}\min\left\{1,\;M^{1-r_{1,i}}\right\}$
@@
=$\overset{(b)}{=}\min\left\{1,\;M^{\Delta-r_{1}}\right\},$
@@
However, Lemma 6.1 is still one step away from the desired bound as it requires $G(S_{11})$ to be connected. Fortunately, as $G(S_{11})$ follows an Erdős–Rényi model, when $p=\Theta\left(\frac{\log n}{n}\right)$, we can show that at least a giant component of size $(1-o(1))\cdot|S_{11}|$ exists in $G(S_{11})$ w.h.p. (see Theorem A.4). In this way, an upper bound similar to (54) can be obtained as follows:
@@
$\mathcal{P}_{\bm{x}^{*}}\left(E_{2,S_{1}}\mid r_{1}\right)\leq n^{-\Theta(\log n)}+\min\left\{1,M^{\Delta-(1-o(1))r_{1}}\right\},$ (55)
@@
The proof is deferred to Appendix B.1. As a result, Lemma 6.2 relaxes the condition of the subgraph $G(S_{11})$ being connected made in Lemma 6.1, and thus can be applied to (51) directly. By symmetry we have the same result for $S_{2}$ as
@@
Plugging these into (51) yields
@@
Now, by plugging (56) into (48) and taking the summation over $r$ and $r^{*}$, we obtain the following upper bound on $\mathcal{P}_{\Delta}$. The details are left in Appendix B.2.
@@
$\mathcal{P}_{\Delta}\leq M^{2\Delta}\cdot\exp\left(-\frac{N_{\Delta}\log n}{n}\left(a+b-(2+o(1))\sqrt{\frac{ab}{M}}\right)\right),$
@@
6.4 Step 4: Sum over $\Delta$
@@
$\leq 2\sum_{\Delta=1}^{\left\lceil n/4\right\rceil}g(\Delta),$ (57)
where $(a)$ uses the property of binomial coefficient that $\binom{n}{k}\leq\left(e\frac{n}{k}\right)^{k}$ and we define each term in the summation as $g(\Delta)$. To proceed, we take the logarithm of $g(\Delta)$ as
@@
where the last step uses $c_{1}\in\left[\frac{2c}{3},c\right]$. This completes the proof.
7 Proof of Theorem 3.2
Our proof technique is mainly inspired by the one of [2, Theorem 1]. Similar to the proof of Theorem 6, we still follow the assumption (38). To begin with, recall the MLE (33) fails to exactly recover the communities $\bm{\kappa}^{*}$ if and only if there exists an alternative hypothesis $\bm{x}=\{\bm{\kappa},\bm{g}\}$ that
@@
7.1 Case 1
We first consider when $\frac{Mp(1-q)}{q(1-p)}\geq 1$, and the MLE fails only if $r\geq r^{*}$. For ease of presentation, without specification, we denote $\mathcal{P}(\cdot)$ as the probability conditioning on the ground truth $\bm{x}^{*}$ and omit the subscript. Then, let us define the following events:
@@ 7.1 Case 1
    node $i\in S_{1}^{*}$ (resp. $S_{2}^{*}$) is connecting to more nodes in $S_{2}^{*}$ (resp. $S_{1}^{*}$) than in $S_{1}^{*}$ (resp. $S_{2}^{*}$), i.e., $|\mathcal{E}(i,S_{1}^{*}\backslash i)|<|\mathcal{E}(i,S_{2}^{*})|$ (resp. $|\mathcal{E}(i,S_{2}^{*}\backslash i)|<|\mathcal{E}(i,S_{1}^{*})|$);
@@
    node $i\in S_{1}^{*}$ (resp. $S_{2}^{*}$) satisfies $g_{ij}=g_{i}$, for any $(i,j)\in\mathcal{E}(i,S_{2}^{*})$ (resp. $\mathcal{E}(i,S_{1}^{*})$).
@@
By symmetry, we have $\mathcal{P}(E_{1})=\mathcal{P}(E_{2})\geq\frac{4}{5}$. Then, consider the event $E=E_{1}\cap E_{2}$ occurs, which indicates there exists a pair of nodes $i\in S_{1}^{*}$ and $j\in S_{2}^{*}$ with the two events $E^{(i)}$ and $E^{(j)}$ occur with some $g_{i}$ and $g_{j}$. As a result, one can construct an alternative hypothesis $\bm{x}$ from $\bm{x}^{*}$ by swapping the community assignment of $i$ and $j$, i.e., $i\in S_{2}$ and $j\in S_{1}$, and let the associated group elements be $g_{i}$ and $g_{j}$, then $\mathcal{P}(\bm{y}|\bm{x})>\mathcal{P}(\bm{y}|\bm{x}^{*})$, indicating $F$ occurs. Therefore,
$\mathcal{P}(F)\geq\mathcal{P}(E_{1}\cap E_{2})\geq\mathcal{P}(E_{1})+\mathcal{P}(E_{2})-1\geq\frac{3}{5},$
@@
$\gamma:=\log^{3}n,\quad\delta:=\frac{\log n}{\log\log n},$
and we assume them to be integers (otherwise round them to the nearest ones). Then, let $H$ be a fixed subset of $S_{1}^{*}$ of size $\frac{n}{\gamma}$, we further define the following two events:
@@
    no node in $H$ connects to at least $\delta$ number of other nodes in $H$,
@@
    node $i\in H$ satisfies $|\mathcal{E}(i,S_{1}^{*}\backslash H)|+\delta\leq|\mathcal{E}(i,S_{2}^{*})|$,
@@
By definition, $E_{H}$ occurs when there exists a node $i\in H$ with some $g_{i}$ that such that $E_{H,A}^{(i)}$ and $E_{B}^{(i,g_{i})}$ occur. Then one can see that $F_{H}\cap E_{H}\subseteq E_{1}$, as $F_{H}\cap E_{H}^{(i)}$ indicates $E^{(i)}$ occurs. Therefore, by applying $\mathcal{P}(F_{H})\geq\frac{9}{10}$ from [2, Lemma 10], we have
@@
$\overset{(b)}{\geq}1-\prod_{i\in H}\left(1-\mathcal{P}\left(E_{H,A}^{(i)}\bigcap E_{B}^{(i,0)}\right)\right),$ (61)
@@
Here, $(a)$ comes from the fact that the two events $E_{H,A}^{(i)}$ and $E_{B}^{(i,0)}$ occurs independently given $r_{1}$ and $r_{2}$; $(b)$ holds as given $r_{1}$ and $r_{2}$, $E_{H,A}^{(i)}$ is deterministic, and by definition $E_{B}^{(i,0)}$ occurs only if $g_{ij}=I$ for all $(i,j)\in\mathcal{E}(i,S_{2}^{*})$, and each $g_{ij}=I$ occurs independent with probability $M^{-1}$. Following this way with some heavy calculation, we obtain
@@
7.2 Case 2
Now we consider $\frac{Mp(1-q)}{q(1-p)}<1$ where MLE fails only if $r\leq r^{*}$. The proof is very similar to Case 1 where most of the result still applies here. Therefore, we will skip those repeated analyses. To begin with, we reuse all the notations of events in Case 1 but redefine the following two: